The length of a wire of a potentiometer is 100 cm and the emf of its standard cell is E volt. It is employed to measure the e.m.f of a battery whose internal resistance is 0.5 ohm. If the balance point is obtained at l = 30 cm from the positive end, the e.m.f of the battery is?

Grade:11

1 Answers

Chetan Naik
6 years ago
By principle of potentiometerE1/E=L1/LSince null point is obtained at 30 cmL1=30L=100 because when E1 is not connected null point is obtained at 100 cmE1/E=30/100E1=30E/100